Please help me solve this equation and explain to me how you did it: 6n-3n=12
Answer by likaaka(51)

well since you are dealing with a single variable, n, you simply combine like terms. 6n-3n=3n
Given that the sum of the original equation is 12, you now know that 3n=12
Now you solve for n by dividing both sides by the coefficient, in this case the 3 so (3n)/3 = 12/3
Your answer is n = 4